OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 5. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (anonymous):

i dont get it

OpenStudy (aum):

Isolate x.

OpenStudy (anonymous):

-5log16 27

OpenStudy (aum):

\(\Large \log_{16}(32) = x + 5 \\ \Large \log_{16}(2*16) = x + 5 \\ \Large \log_{16}(2) + \log_{16}(16) = x + 5 \\ \Large \log_{16}(2) + 1 = x + 5 \\ \Large \text{Subtract 5:} \\ \Large \log_{16}(2) -4 = x \\ \Large x = \log_{16}(2) -4 \\ \Large x = \frac{1}{\log_{2}(16)} -4 \\ \Large x = \frac 14 -4 \\ \Large x = \frac{-15}{4} \\ \Large x = -3\frac 34 \)

OpenStudy (anonymous):

damm

OpenStudy (aum):

A quicker way to do this problem may be to use the change of base formula: \(\Large \log_{16}32 = \frac{\log_{2}32}{\log_216} = \frac 54 \)

OpenStudy (anonymous):

oh

OpenStudy (aum):

So the left hand side is 5/4 5/4 = x + 5 x = 5/4 - 5 = (5-20)/4 = -15/4 = -3 & 3/4

OpenStudy (anonymous):

so the answer is -3 3/4 or 5/4

OpenStudy (aum):

No the answer is -3 3/4 solving by two different methods.

OpenStudy (anonymous):

ohhh